# Generalized Chern-Pontryagin models

### v2

@@ Abstract
We formulate a new class of modified gravity models, that is, generalized four-dimensional Chern-Pontryagin models, whose action is characterized by an arbitrary function of the Ricci scalar $R$ and the Chern-Pontryagin topological term ${}^{*}RR$, i.e., $f(R,{}^{*}RR)$. Within this framework, we derive the gravitational field equations and solve them for a particular model, $f(R,{}^{*}RR)=R+\beta({}^{*}RR)^{2}$, considering two ansatzes: the slowly rotating metric and first-order perturbations of Gödel-type metrics. For the former, we find a first-order correction to the frame-dragging effect boosted by the parameter $L$, which characterizes the departures from general relativity results. For the latter, Gödel-type metrics hold unperturbed. We conclude this paper by displaying that generalized four-dimensional Chern-Pontryagin models admit a scalar-tensor representation, whose explicit form presents two scalar fields: $\Phi$, a dynamical degree of freedom, while the second, $\vartheta$, a non-dynamical degree of freedom. In particular, the scalar field $\vartheta$ emerges coupled with the Chern-Pontryagin topological term ${}^{*}RR$, i.e., $\vartheta{}^{*}RR$, which is nothing more than Chern-Simons term.
@@ I Introduction
Studies of gravitational phenomena are one of the most important research lines of the contemporary physics. Such an interest is inspired in part by important recent discoveries, as late-time accelerating expansion of the Universe Riess , detection of gravitational waves LIGO and obtaining an image of the shadow of a black hole Akiyama . Within this context, modified gravity models have played a pivotal role since they can tackle with the open questions in gravitational physics or, in the cosmological scenario, provide alternative answers to $\Lambda$CDM without making use of the addition of a certain amount of unseen exotic matter/energy Clifton ; Ourbook ; Olmo2011 .
@@
As it is known, various extensions of Einstein’s gravity have been proposed – those including modification in the pure gravity sector Sot ; DeFelice ; Noj ; Nojiri:2017ncd and others involving extra fields, either scalar Brans ; Horndeski ; Koba , vector KosGra ; Jacob ; Del or even tensor ones (that is, so-called bimetric gravity). One of such extensions, that has attracted a strong interest in recent years, is the four-dimensional Chern-Simons modified gravity (CSMG) originally proposed in JaPi . In essence, the CSMG is the theory whose Lagrangian includes, besides of the scalar curvature, the Chern-Pontryagin topological term ${}^{*}RR\equiv\frac{1}{2}\epsilon^{\mu\nu\lambda\rho}R_{\mu\nu\alpha\beta}R_{\lambda\rho}^{\phantom{\lambda\rho}\alpha\beta}$, multiplied by an extra scalar $\vartheta$ called the Chern-Simons (CS) coefficient, i.e., the $\vartheta{}^{*}RR$ term (also called the CS term) is added. This term, being the natural four-dimensional extension of the 3-dimensional gravitational Chern-Simons term DJT , displays numerous interesting properties. First, it can break parity explicitly, by depending on the form of $\vartheta$; if $\vartheta$ is even under parity transformation, once ${}^{*}RR$ is odd under parity transformation. Also, it can break Lorentz symmetry at the action level, since one can rewrite the CS term as an interaction term between a background vector $v_{\mu}\equiv\partial_{\mu}\vartheta$ and the topological Chern-Simons current $K^{\mu}$ (for a discussion of other Lorentz/parity-breaking gravity models see KosGra ; Li ; Li2 ; Li3 ; Li4 ; Rao ). Second, it was checked that CSMG shares some solutions with the general relativity (GR), for example, the Schwarzschild metric. Conversely, one can mention, as a counterexample, the Kerr metric which is not a solution of CSMG, as discussed in Yunes ; Konno ; Alexander:2009tp . In the further generalization of the CSMG, the so-called dynamical CSMG (DCSMG), the kinetic term for $\vartheta$ is introduced. A detailed discussion of the consistency of various known metrics within the DCSMG is presented in Grumiller , and various issues related to gravitational waves in metric DCSMG were studied in Nojiri:2019nar ; Nojiri:2020pqr . A metric-affine extension of the DCSMG with projective invariance has been developed Boudet , and propagation of gravitational waves within this theory has also been investigated Boudet2 ; Bombacigno . The Chern-Pontryagin topological invariant also appears naturally in the context of string/M-theory, specifically, compactifying $D=11$ M-theory down to $D=4$ effective gravity theories. In particular, one can cite the Starobinsky-Bel-Robinson (SBR) model Ketov:2022lhx ; Ketov:2022zhp , whose action involves quadratic Chern-Pontryagin terms, $({}^{*}RR)^{2}$, its cosmological aspects were studied in Ketov:2022lhx ; Ketov:2022zhp . Another interesting string-inspired model closely related to SBR model is the Einstein-Grisaru-Zanon effective gravity theory which has been investigated in CamposDelgado:2024jst .
@@
In this paper, driven by $f(R)$ theories, we propose a more generic class of theories called generalized Chern-Pontryagin models, whose action, instead of being linear or quadratic in ${}^{*}RR$, displays a generic functional dependency on this object and also the Ricci scalar. Certainly, this will allow for a richer framework to explore, among other issues, new parity-violating solutions of the modified field equations in the model. At the same time, many well-known metrics, including Schwarzschild and Gödel ones, actually, all metrics fulfilling the Pontryagin constraint ${}^{*}RR=0$, can be easily shown to be simultaneous solutions of GR and generalized Chern-Pontryagin theories. Then, it is interesting to examine the influence of the generalized Chern-Pontryagin term under small perturbations of these metrics within the modified field equations. Nonetheless, our main goal is to provide a natural mechanism to generate the four-dimensional CS term, which was originally proposed in analogy to the electromagnetism JaPi .
@@
The structure of the paper looks like follows. In Section II, we write down the classical action of the theory and obtain the corresponding equations of motion. In Section III, we study the consistency of slowly rotating black hole solutions and first-order perturbations of ST-homogeneous Gödel-type metrics in the modified theory. In Section IV, we provide a mechanism to generate the Chern-Simons term. Finally, in Section IV we discuss our results.
@@ II Generalized Chern-Pontryagin gravity models
Note that the modified field equations for $f_{2}(^{*}RR)\propto\left(\,{}^{*}RR\right)^{n}$, with $n\geq 1$, reduce to the GR ones for a variety of metric classes, for example, static spherically symmetric, Friedmann-Robertson-Walker, (A)dS and Gödel-type metrics, since ${}^{*}RR=0$, for all these metrics (actually, this condition is nothing more than the Pontryagin constraint satisfied for certain metrics once the specific relation between corresponding Newman-Penrose scalars is fulfilled, see the discussion in Grumiller ). In principle, this constraint holds for some other metrics as well. However, notice that to yield non-trivially modified field equations of motion it is necessary to have $C_{\mu\nu}\neq 0$ which leads to $\,{}^{*}RR\neq 0$ in particular cases. For example, just this situation occurs for the Kerr metric, as shown in Yunes . We note also that had we considered the “agravity-like” theory which does not involve Einstein-Hilbert term (cf. Salvio ), so that $f_{1}(R)=0$, and only higher derivatives are present in the action, the metrics satisfying the Pontryagin constraint would solve the equations of motion in a trivial manner since in this case $f_{R}=0$ immediately, and $f_{\,{}^{*}RR}=0$ if $f({}^{*}RR)\propto({}^{*}RR)^{n}$, with $n\geq 1$. In the following section, we shall investigate perturbations of some metrics within the generalized Chern-Pontryagin gravity model.
@@ III.1 Perturbation of the Schwarzschild spacetime
where $M$ is the mass of any spherically symmetric mass distribution and the function $\omega(r,\theta)$ is of the first order in $\epsilon\sim J/M^{2}$, here $J$ is angular momentum. In this perturbation scheme, $\epsilon$ is considered to be a small parameter. Hereafter, we take account of equations up to the first order in $\epsilon$. In this situation, the Chern-Pontryagin term for the metric (9) up to first order in $\epsilon$ is non-zero
@@ III.2 Perturbative scheme for ST-homogeneous Gödel-type metrics
The original Gödel solution Godel represents the particular case $m^{2}=2\omega^{2}$, then it is a particular example of the hyperbolic class of ST-homogeneous Gödel-type metrics. For brevity, from now on, we shall refer to ST-homogeneous Gödel-type metrics simply as to Gödel-type metrics. It was shown in RT that all three classes of these metrics present at least one region ($r_{1}<r<r_{2}$), where there exists CTCs defined by the circle $C=\{(t,r,\theta,z);t=t_{0},r=r_{0},\theta\in[0,2\pi],z=z_{0}\}$. In the especial case of the hyperbolic class corresponding to $m^{2}\geq 4\omega^{2}$, the presence of CTCs is completely avoidable. It can be seen by using the critical radius  [^2] The critical radius is defined as the maximum radius allowing for the presence of CTCs. ($r_{c}$) equation given by
@@
Notice that for $m^{2}\geq 4\omega^{2}$, the above equation do not provide any real $r_{c}$. Earlier, the consistency of the three classes of Gödel-type metrics within various modified gravity models have been discussed in a set of papers (see e.g. ourReb for the Horava-Lifshitz case, and BD for the Brans-Dicke (BD) case). In particular, it was shown in Godelour1 ; Godelour2 ; Godelour3 that ${}^{*}RR=0$ for Gödel-type metrics, as a consequence the contribution of $C$-tensor for the field equations (2) vanishes. In this subsection, we shall examine the possibility of breaking the ST-homogeneity for Gödel-type metrics within the model $f({}^{*}RR)=R+\beta(^{*}RR)^{2}$. To do that, we shall introduce a perturbative approach, which will be discussed in detail below. Then we can straightforwardly substitute the perturbed metric into the field equations to find the full solutions up to first order in the perturbation parameter.
@@
Now, we solve the modified Einstein equation (2) for the particular model $f({}^{*}RR)=R+\beta(^{*}RR)^{2}$. To start with, let us consider the $(t,z)$-component
@@
$(1+\xi\,h_{2}(r,\theta,z))^{6}\left(8\,\frac{\partial}{\partial\theta}h_{1}(r,\theta,z)-\frac{\partial}{\partial\theta}h_{2}(r,\theta,z)\right)+$ (35)
@@
Requiring $m^{2}\neq 4\omega^{2}$, we find that the only solution is the trivial one, $h_{1}(r,\theta,z)=h_{2}(r,\theta,z)=0$. Thus, one concludes that the particular generalized Pontryagin model $f(R,{}^{*}RR)=R+\beta(^{*}RR)^{2}$ does not support first-order perturbations of Gödel-type metrics, unlike the non-dynamical Chern-Simons modified gravity, in which the non-trivial first-order contributions were found Alt .

### v3

@@ Abstract
We formulate a new class of modified gravity models, that is, generalized four-dimensional Chern-Pontryagin models, whose action is characterized by an arbitrary function of the Ricci scalar $R$ and the Chern-Pontryagin topological term ${}^{*}RR$, i.e., $f(R,{}^{*}RR)$. Within this framework, we derive the gravitational field equations and solve them for the particular models, $f(R,{}^{*}RR)=R+\beta({}^{*}RR)^{2}$ and $f(R,{}^{*}RR)=R+\alpha R^{2}+\beta({}^{*}RR)^{2}$, considering two ansatzes: the slowly rotating metric and first-order perturbations of Gödel-type metrics. For the former, we find a first-order correction to the frame-dragging effect boosted by the parameter $L$, which characterizes the departures from general relativity results. For the latter, Gödel-type metrics hold unperturbed, for specific sort of perturbed metric functions. We conclude this paper by displaying that generalized four-dimensional Chern-Pontryagin models admit a scalar-tensor representation, whose explicit form presents two scalar fields: $\Phi$, a dynamical degree of freedom, while the second, $\vartheta$, a non-dynamical degree of freedom. In particular, the scalar field $\vartheta$ emerges coupled with the Chern-Pontryagin topological term ${}^{*}RR$, i.e., $\vartheta{}^{*}RR$, which is nothing more than Chern-Simons term.
@@ I Introduction
Studies of gravitational phenomena are one of the most important research lines in contemporary physics. Such an interest is inspired in part by important recent discoveries, such as the late-time accelerating expansion of the Universe Riess , detection of gravitational waves LIGO and obtaining an image of the shadow of a black hole Akiyama . Within this context, modified gravity models have played a pivotal role since they can tackle open questions in gravitational physics or, in the cosmological scenario, provide alternative answers to $\Lambda$CDM without making use of the addition of a certain amount of unseen exotic matter/energy Clifton ; Ourbook ; Olmo2011 .
@@
As it is known, various extensions of Einstein’s gravity have been proposed – those including modification in the pure gravity sector Sot ; DeFelice ; Noj ; Nojiri:2017ncd and others involving extra fields, either scalar Brans ; Horndeski ; Koba , vector KosGra ; Jacob ; Del or even tensor ones (that is, so-called bimetric gravity). One of such extensions, that has attracted a strong interest in recent years, is the four-dimensional Chern-Simons modified gravity (CSMG) originally proposed in JaPi . In essence, the CSMG is the theory whose Lagrangian includes, besides of the scalar curvature, the Chern-Pontryagin topological term ${}^{*}RR\equiv\frac{1}{2}\epsilon^{\mu\nu\lambda\rho}R_{\mu\nu\alpha\beta}R_{\lambda\rho}^{\phantom{\lambda\rho}\alpha\beta}$, multiplied by an extra scalar $\vartheta$ called the Chern-Simons (CS) coefficient, i.e., the $\vartheta{}^{*}RR$ term (also called the CS term) is added. This term, being the natural four-dimensional extension of the 3-dimensional gravitational Chern-Simons term DJT , displays numerous interesting properties. First, it can break parity explicitly, depending on the form of $\vartheta$; if $\vartheta$ is even under parity transformation, once ${}^{*}RR$ is odd under parity transformation. Also, it can break Lorentz symmetry at the action level, since one can rewrite the CS term as an interaction term between a background vector $v_{\mu}\equiv\partial_{\mu}\vartheta$ and the topological Chern-Simons current $K^{\mu}$. Second, it was checked that CSMG shares some solutions with the general relativity (GR), for example, the Schwarzschild metric. Conversely, as a counterexample, one can mention the Kerr metric which is not a solution of CSMG, as discussed in Yunes ; Konno ; Alexander:2009tp . In the further generalization of the CSMG, the so-called dynamical CSMG (DCSMG), the kinetic term for $\vartheta$ is introduced.
A detailed discussion of the consistency of various known metrics within the DCSMG is presented in Grumiller , and various issues related to gravitational waves in metric DCSMG were studied in Nojiri:2019nar ; Nojiri:2020pqr . A metric-affine extension of the DCSMG with projective invariance has been developed Boudet , and propagation of gravitational waves within this theory has also been investigated Boudet2 ; Bombacigno . The Chern-Pontryagin topological invariant also appears naturally in the context of string/M-theory, specifically, compactifying $D=11$ M-theory down to $D=4$ effective gravity theories. In particular, one can cite the Starobinsky-Bel-Robinson (SBR) model Ketov:2022lhx ; Ketov:2022zhp , whose action involves quadratic Chern-Pontryagin terms, $({}^{*}RR)^{2}$, its cosmological aspects were studied in Ketov:2022lhx ; Ketov:2022zhp . Another interesting string-inspired model closely related to SBR model is the Einstein-Grisaru-Zanon effective gravity theory which has been investigated in CamposDelgado:2024jst .
One more motivation for studying the modified gravity theories, especially those ones involving the ${}^{*}RR$ term and its generalizations, is related with the possibility of breaking of Lorentz and/or CPT symmetries suggested already in the seminal papers KosSam , where this possibility has been justified by string theory (further, other motivations for the possibility of Lorentz-CPT symmetry breaking, based on loop quantum gravity, minimal length hypothesis, space-time noncommutativity, etc. were introduced), therefore, various Lorentz-CPT breaking extensions of gravity can be defined (for a review of such extensions, see KosLi ). The simplest known example of the Lorentz-CPT breaking term in gravity is just the four-dimensional gravitational CS term JaPi . Already in KosGra possibilities to generalize this term, and hence to obtain higher-order Lorentz-CPT breaking terms in gravity were mentioned. Therefore, the natural idea consists in studying of gravity models involving additive terms of the form $({}^{*}RR)^{n}$ which, first, generalize the gravitational CS term, second, break the CPT symmetry for odd values of $n$, representing the most natural Riemannian CPT-breaking gravity models (we note that there are also non-Riemannian CPT-breaking theories Li ; Porf1 ; Li2 ; Porf2 ; Li3 ; Li4 ; Porf3 ; Porf4 ; Rao ; Porf5 ; Porf6 ).
@@
In this paper, driven by $f(R)$ theories, we propose a more generic class of theories called generalized Chern-Pontryagin models, whose action, instead of being linear or quadratic in ${}^{*}RR$, displays a generic functional dependency on this object and also the Ricci scalar. Certainly, this will allow for a richer framework to explore, among other issues, new parity-violating solutions of the modified field equations in the model. At the same time, many well-known metrics, including Schwarzschild and Gödel ones, actually, all metrics fulfilling the Pontryagin constraint ${}^{*}RR=0$, can be easily shown to be simultaneous solutions of GR and generalized Chern-Pontryagin theories. Then, it is interesting to examine the influence of the generalized Chern-Pontryagin term under small perturbations of these metrics within the modified field equations. Nonetheless, our main goal is to provide a natural mechanism to generate the four-dimensional CS term, which was originally proposed in analogy to electrodynamics JaPi .
@@
The structure of the paper looks like follows. In Section II, we write down the classical action of the theory and obtain the corresponding equations of motion. In Section III, we study the consistency of slowly rotating black hole solutions and first-order perturbations of ST-homogeneous Gödel-type metrics in the modified theory. In Section IV, we provide a mechanism to generate the Chern-Simons term. Finally, in Section V we discuss our results.
@@ II Generalized Chern-Pontryagin gravity models
Note that the modified field equations for $f_{2}(^{*}RR)\propto\left(\,{}^{*}RR\right)^{n}$, with $n\geq 1$, reduce to the GR ones for a variety of metric classes, for example, static spherically symmetric, Friedmann-Robertson-Walker, (A)dS and Gödel-type metrics, since ${}^{*}RR=0$, for all these metrics (actually, this condition is nothing more than the Pontryagin constraint satisfied for certain metrics once the specific relation between corresponding Newman-Penrose scalars is fulfilled, see the discussion in Grumiller ). In principle, this constraint holds for some other metrics as well. However, notice that to yield non-trivially modified field equations of motion it is necessary to have $C_{\mu\nu}\neq 0$ which leads to $\,{}^{*}RR\neq 0$ in particular cases. For example, just this situation occurs for the Kerr metric, as shown in Yunes . We note also that had we considered the “agravity-like” theory which does not involve Einstein-Hilbert term (cf. Salvio ), so that $f_{1}(R)=0$, and only higher derivatives are present in the action, the metrics satisfying the Pontryagin constraint would solve the equations of motion in a trivial manner since in this case $f_{R}=0$ immediately, and $f_{\,{}^{*}RR}=0$ if $f({}^{*}RR)\propto({}^{*}RR)^{n}$, with $n>1$. In the following section, we shall investigate perturbations of some metrics within the generalized Chern-Pontryagin gravity model.
@@ III.1 Perturbation of the Schwarzschild spacetime
where $M$ is the mass of any spherically symmetric mass distribution and the function $\omega(r,\theta)$ is of the first order in $\epsilon\sim J/M^{2}$, here $J$ is angular momentum. In this perturbation scheme, $\epsilon$ is considered to be a small parameter. Hereafter, we take account of equations up to the first order in $\epsilon$. In this situation, the Chern-Pontryagin term for the metric (9) up to first order in $\epsilon$ is non-zero:
@@ III.2 Perturbative scheme for ST-homogeneous Gödel-type metrics
The original Gödel solution Godel represents the particular case $m^{2}=2\omega^{2}$, then it is a particular example of the hyperbolic class of ST-homogeneous Gödel-type metrics. For brevity, from now on, we shall refer to ST-homogeneous Gödel-type metrics simply as to Gödel-type metrics. It was shown in RT that all three classes of these metrics present at least one region ($r_{1}<r<r_{2}$), where there exist CTCs defined by the circle $C=\{(t,r,\theta,z);t=t_{0},r=r_{0},\theta\in[0,2\pi],z=z_{0}\}$. In the special case of the hyperbolic class corresponding to $m^{2}\geq 4\omega^{2}$, the presence of CTCs is completely avoidable. It can be seen by using the critical radius  [^2] The critical radius is defined as the maximum radius allowing for the presence of CTCs. ($r_{c}$) equation given by
@@
Notice that for $m^{2}\geq 4\omega^{2}$, the above equation does not provide any real $r_{c}$. Earlier, the consistency of the three classes of Gödel-type metrics within various modified gravity models have been discussed in a set of papers (see e.g. ourReb for the Horava-Lifshitz case, and BD for the Brans-Dicke (BD) case). In particular, it was shown in Godelour1 ; Godelour2 ; Godelour3 that ${}^{*}RR=0$ for Gödel-type metrics, as a consequence the contribution of $C$-tensor for the field equations (2) vanishes. In this subsection, we shall examine the possibility of breaking the ST-homogeneity for Gödel-type metrics within the model $f(R,{}^{*}RR)=R+\beta(^{*}RR)^{2}$. To do that, we shall introduce a perturbative approach, which will be discussed in detail below. Then we can straightforwardly substitute the perturbed metric into the field equations to find the full solutions up to first order in the perturbation parameter.
@@
Now, we solve the first-order perturbed modified Einstein equation (2) for the particular model $f({}^{*}RR)=R+\beta(^{*}RR)^{2}$. To start with, let us consider the $(t,z)$-component
@@
$\left(8\,\frac{\partial}{\partial\theta}h_{1}(r,\theta,z)-\frac{\partial}{\partial\theta}h_{2}(r,\theta,z)\right)+$ (35)
@@
Requiring $m^{2}\neq 4\omega^{2}$, we find that the only solution is the trivial one, $h_{1}(r,\theta,z)=h_{2}(r,\theta,z)=h_{3}(r,\theta,z)=0$.
It should be realized that Eqs. (33), (34) and (35) are automatically satisfied by considering particular metric perturbations that depend only on either the $(r,\theta)$ or $(r,z)$ coordinates. The former case  [^3] Physically speaking, this case means that the metric perturbations preserve the axial symmetry along the $z$-direction., $h_{i}(r,\theta,z)=h_{i}(r,\theta)$, is trivial since it leads to the vanishing of the Chern-Pontryagin term (29), which in turn results in the vanishing of the $C$-tensor, thereby enforcing the field equations to reduce to those of GR. As for the latter case  [^4] In this case, the metric perturbations break the axial symmetry along the $z$-direction., $h_{i}(r,\theta,z)=h_{i}(r,z)$, it requires a more careful inspection of the remaining field equations. In particular, the $(r,z)-$component of the modified Einstein equations is given by
$\frac{\partial}{\partial z}h_{2}(r,z)\frac{dD(r)}{dr}-\frac{\partial}{\partial z}h_{3}(r,z)\frac{dD(r)}{dr}+2\omega H(r)\frac{\partial}{\partial z}h_{0}(r,z)-2\omega H(r)\frac{\partial}{\partial z}h_{1}(r,z)+$ (36)
$+$ $0\displaystyle 2D(r)\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)+D(r)\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)=$ 0 .
A particular solution is found by imposing the following relationships among perturbed metric functions, $h_{0}(r,z)=h_{1}(r,z),h_{2}(r,z)=h_{3}(r,z)=-2h_{1}(r,z)$. The $zz-$component
$2h_{1}(r,z)(m^{2}-3\omega^{2})=0,$ (37)
is an algebraic equation which the only non-trivial solution is $h_{1}(r,z)=0$.
@@
Thus, one concludes that the particular generalized Pontryagin model $f(R,{}^{*}RR)=R+\beta(^{*}RR)^{2}$ does not support first-order perturbations of Gödel-type metrics discussed above. In contrast, the Chern-Simons modified gravity exhibits non-trivial first-order contributions Alt .
Another important example to consider is a generalization of the Starobinsky model, which corresponds to taking $f(R,{}^{*}RR)=R+\alpha R^{2}+\beta(^{*}RR)^{2}$. In this situation, we proceed similarly to the previous case to solve the field equations. Nevertheless, we impose a simplification, as the field equations become cumbersome to tackle due to the presence of higher-order derivative terms. In particular, we consider only perturbed metric functions that do not explicitly depend on the coordinate $\theta$, that is, $h_{i}(r,\theta,z)=h_{i}(r,z)$. Recalling that the choice $h_{i}(r,\theta,z)=h_{i}(r,\theta)$ is not well-motivated because Chern-Pontryagin corrections vanish, as explained before. To begin with, let us start by considering the $(r,z)-$component of the field equations
$\Bigg{[}-2D(r)\left(\frac{\partial^{4}}{\partial z\partial r^{3}}h_{2}(r,z)\right)-4D(r)\left(\frac{\partial^{4}}{\partial z\partial r^{3}}h_{0}(r,z)\right)-4\left(\frac{\partial^{3}}{\partial z\partial r^{2}}h_{2}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$-4\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)\left(\frac{\mathrm{d}^{2}}{\mathrm{\leavevmode\nobreak\ d}r^{2}}D(r)\right)-4\left(\frac{\partial^{3}}{\partial z\partial r^{2}}h_{0}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$-2D(r)\left(\frac{\partial^{4}}{\partial z^{3}\partial r}h_{2}(r,z)\right)+2\left(\frac{\partial^{2}}{\partial z\partial r}h_{3}(r,z)\right)\left(\frac{\mathrm{d}^{2}}{\mathrm{\leavevmode\nobreak\ d}r^{2}}D(r)\right)+2\left(\frac{\partial^{3}}{\partial z\partial r^{2}}h_{3}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$-4D(r)\left(\frac{\partial^{4}}{\partial z^{3}\partial r}h_{0}(r,z)\right)-2D(r)\left(\frac{\partial^{4}}{\partial z^{3}\partial r}h_{3}(r,z)\right)$
$-4\left(\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)\right)\left(\frac{\mathrm{d}^{2}}{\mathrm{\leavevmode\nobreak\ d}r^{2}}D(r)\right)+4D(r)\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}\ln{D(r)}\right)^{2}$
$-4\omega H(r)\left(\frac{\partial^{2}}{\partial z\partial r}h_{1}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}\ln{D(r)}\right)+8D(r)\omega^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{1}(r,z)\right)$
$-4\left(\frac{\mathrm{\leavevmode\nobreak\ d}}{\mathrm{\leavevmode\nobreak\ d}r}H(r)\right)\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)\omega-2m^{2}\left(\frac{\partial}{\partial z}h_{2}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$-4\omega^{3}H(r)\left(\frac{\partial}{\partial z}h_{1}(r,z)\right)+28D(r)\omega^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)-2\omega^{2}\left(\frac{\partial}{\partial z}h_{3}(r,z\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$+4\omega^{3}H(r)\left(\frac{\partial}{\partial z}h_{0}(r,z)\right)-4H(r)\left(\frac{\partial^{3}}{\partial z\partial r^{2}}h_{0}(r,z)\right)\omega$
$+2m^{2}\left(\frac{\partial}{\partial z}h_{3}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)-14D(r)m^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)\right)+2\omega^{2}\left(\frac{\partial}{\partial z}h_{2}(r,z\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$+4H(r)\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)\omega\left(\frac{\mathrm{\leavevmode\nobreak\ d}}{\mathrm{\leavevmode\nobreak\ d}r}\ln{D(r)}\right)-4m^{2}H(r)\omega\left(\frac{\partial}{\partial z}h_{0}(r,z)\right)$
$+4m^{2}H(r)\left(\frac{\partial}{\partial z}h_{1}(r,z)\right)\omega-2\left(\frac{\partial^{2}}{\partial z\partial r}h_{3}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}\ln{D(r)}\right)^{2}D(r)$
$+4\left(\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}\ln{D(r)}\right)^{2}D(r)+10D(r)\omega^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)\right)$
$-8D(r)m^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{3}(r,z)\right)+8D(r)\left(\frac{\partial^{2}}{\partial z\partial r}h_{3}(r,z)\right)\omega^{2}$
$+4\omega\left(\frac{\mathrm{\leavevmode\nobreak\ d}}{\mathrm{\leavevmode\nobreak\ d}r}H(r)\right)\left(\frac{\partial^{2}}{\partial z\partial r}h_{1}(r,z)\right)-28D(r)m^{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)$
$+4\omega H(r)\left(\frac{\partial^{3}}{\partial z\partial r^{2}}h_{1}(r,z)\right)\Bigg{]}\alpha+H(r)\left(\frac{\partial}{\partial z}h_{1}(r,z)\right)\omega+\frac{1}{2}\left(\frac{\partial}{\partial z}h_{3}(r,z)(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)$
$-\frac{1}{2}\left(\frac{\partial}{\partial z}h_{2}(r,z)\right)\left(\frac{\mathrm{d}}{\mathrm{\leavevmode\nobreak\ d}r}D(r)\right)-\left(\frac{\partial^{2}}{\partial z\partial r}h_{0}(r,z)\right)D(r)$
$-H(r)\left(\frac{\partial}{\partial z}h_{0}(r,z)\right)\omega-\frac{1}{2}\left(\frac{\partial^{2}}{\partial z\partial r}h_{2}(r,z)\right)D(r)=0.$ (38)
Note that this is a high-order derivative partial differential equation (PDE). The particular relationship among the perturbed metric functions, $h_{0}(r,z)=h_{1}(r,z),h_{2}(r,z)=h_{3}(r,z)=-2h_{1}(r,z)$ and $h_{1}(r,z)=F(r)+G(z)$, solves Eq.(38). Although we have obtained a simple solution for the $(r,z)-$component, this solution does not satisfy the other components of the field equations. Indeed, the more natural conclusion is that the field equations present a highly overconstrained set of PDEs, whose only possible solution is the trivial one., i.e, $h_{0}(r,z)=h_{1}(r,z)=h_{2}(r,z)=h_{3}(r,z)=0$.
Therefore, one concludes that by imposing the axial symmetry breaking along the $z$-direction, $f(R,{}^{*}RR)=R+\alpha R^{2}+\beta(^{*}RR)^{2}$ does not support first-order perturbations of Gödel-type metrics.